# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ 3.2. Definition of Greenwashing
As climate-related disclosures face increasing regulation and greenwashing poses significant risks to a company’s reputation, some companies adopt a strategy of silence, avoiding discussions about their environmental impact. This deliberate lack of communication is known as greenhushing (Letzing, 2022). These definitions imply that greenwashing is a deliberate act; however, in many cases, it results from an error or miscommunication by companies genuinely trying to showcase their sustainability efforts, and that are trying to best follow disclosure standards.
As shown by Definitions 3.1 and 3.2, greenwashing is not defined by easily identifiable properties but as a general concept. Since the concept is so unspecific, researchers focused on components indicative of potentially misleading communications but easier to define. Because of this, in this review we are mentioning “greenwashing” explicitly, but also paraphrasing it as “misleading communications”, “misrepresentation of the company’s environmental impact/stance/performance”, or mentioning only components of it such as “cheap talk”, “selective disclosure/transparency”, “deceptive techniques”, “biased narrative”. They should all be understood in the context of climate-related misleading communications, as components associated or indicative of potential greenwashing even if they are not synonymous.
4. Intermediary tasks
4.1. Pretraining Models on Climate-Related Text
The first step in applying a language model for a given task is typically the pretraining process, which involves training the model on relevant corpora. Although general-purpose models such as BERT (Devlin et al., 2019) and LLaMA (Touvron et al., 2023) have demonstrated strong performance in various tasks, they are inherently limited by the knowledge and vocabulary present in their training corpora. This limitation can lead to suboptimal results when these models are applied to highly specialized domains with unique terminology and concepts. Hence, domain-specific models are usually trained on targeted datasets to improve their performance in niche areas.
@@ 4.1. Pretraining Models on Climate-Related Text
A language model is pretrained on climate-related text, usually on the classical tasks of masked token prediction and next sentence prediction. The goal is to produce a domain-specific pre-trained language model that will be subsequently fine-tuned to specific tasks. There are two approaches: (1) training a model with a domain-specific corpus from scratch, or (2) further-training [^2] Further-training refers to the process of training a pretrained model on its original task (e.g., next-token prediction or masked-token prediction) using additional domain-specific data to specialize its knowledge for that domain. a generalist pretrained model on a domain-specific corpus.
@@
There have been several efforts to create a climate-related corpus. Webersinke et al. (2022) gathered climate-related news articles, research abstracts, and corporate reports. Vaghefi et al. (2022) introduced Deep Climate Change, a dataset composed of abstracts of articles from climate scientists, and built a corpus specific to climate research texts. Schimanski et al. (2023b) introduced a dataset that focuses on text related to Environment Social and Governance (ESG). Similarly, Mehra et al. (2022) built a dataset using text from the Knowledge Hub of Accounting for Sustainability for an ESG domain-specific corpus. More recently, Thulke et al. (2024) proposed a dataset comprising news, publications (abstracts and articles), books, patents, the English Wikipedia, policy and finance-related texts, Environmental Protection Agency documents, and ESG, and IPCC reports. Mullappilly et al. (2023) proposed a climate-specific multilingual dataset. Yu et al. (2024) published a pretraining dataset constructed with annual and sustainable reports from EU banks. Unfortunately, to the best of our knowledge, only Yu et al. (2024)’s ClimateBUG-data dataset is publicly available.
Supervised Solutions
Based on the above datasets, the models ClimateBERT, ClimateGPT-2, climateBUG-LM and EnvironmentalBERT, ESGBERT were proposed (Webersinke et al., 2022; Vaghefi et al., 2022; Yu et al., 2024; Schimanski et al., 2023b; Mehra et al., 2022). More recently, generative domain-specific models such as Llama-2 for ClimateGPT (Thulke et al., 2024) or Vicuna for Arabic Mini-ClimateGPT (Mullappilly et al., 2023) have also been proposed.
@@
The domain-specific models drastically improve the performance on domain-specific masked-language modeling (Webersinke et al., 2022; Yu et al., 2024) and next token prediction (Vaghefi et al., 2022). They were also evaluated on domain-specific downstream tasks. These tasks are either based on pre-existing datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) or introduced by the authors, as in ClimateBERT’s climate detection (Webersinke et al., 2022). At this stage, we evaluate whether fine-tuning enhances downstream performance, with a detailed analysis of the tasks presented in subsequent sections.
Cl.BERT DisRoBERTa NB
Climatext (Spokoyny et al., 2023; Leippold and Varini, 2020) 85.14 86.06 83.39
Climate Detection (Webersinke et al., 2022) 99.1±$\pm 1$ 98.6±$\pm 0.8$
Sentiment (Webersinke et al., 2022) 83.8±$\pm 3.6$ 82.5±$\pm 4.6$
Net-zero/Reduction (Schimanski et al., 2023a) 96.2±$\pm 0.4$ 94.4±$\pm 0.6$
Cl.BERT DisRoBERTa TF-IDF
TCFD classification (Sampson et al., 2022) 0.852 0.819 0.867
Cl.BERT DisRoBERTa SVM
SciDCC (Spokoyny et al., 2023; Mishra and Mittal, 2021) 52.97 51.13 48.02
ClimaTOPIC (Spokoyny et al., 2023) 64.24 63.61 58.34
Cl.FEVER (claim) (Xiang and Fujii, 2023; Leippold and Diggelmann, 2020) 76.8 72
Cl.FEVER (claim) (Webersinke et al., 2022; Leippold and Diggelmann, 2020) 75.7±$\pm 4.4$ 74.8±$\pm 3.6$
Cl.FEVER (evid.) (Spokoyny et al., 2023; Leippold and Diggelmann, 2020) 61.54 61.54
Cl.Stance (Spokoyny et al., 2023; Vaid et al., 2022) 52.84 52.51 42.92
ClimateEng (Spokoyny et al., 2023; Vaid et al., 2022) 71.83 72.33 51.81
ClimaINS (Spokoyny et al., 2023) 84.80 84.38 86.00
ClimaBENCH (Spokoyny et al., 2023) 69.44 69.27
Nature (Schimanski et al., 2024a) 94.11 94.03
@@
Commitment&Actions (Bingler et al., 2023) 81 79 76
Specificity (Bingler et al., 2023) 77 76 75
Cl.BERT DisBERT SVM
Env. Claim (Stammbach et al., 2023) 83.8 83.7 70.9
Cl.BERT DisBERT LSTM
DARE sentiment (Xiang and Fujii, 2023) 87.4 89.9 88.2
(a) F1-scores of climateBERT when compared to a similarly sized models. ROC-AUC for TCFD classification.
@@
Environment (Schimanski et al., 2023b) 92.35±$\pm 2.29$ 93.19±$\pm 1.65$ 90.97±$\pm 2$ $92.35\pm 1.65$
Social (Schimanski et al., 2023b) 89.87±$\pm 1.35$ 91.90±$\pm 1.79$ 90.59±$\pm 1.03$ $91.24\pm 1.86$
Governance (Schimanski et al., 2023b) 77.03±$\pm 1.82$ 78.48±$\pm 2.62$ 76.65±$\pm 2.39$ 78.86±$\pm 1.59$
@@
Water (Schimanski et al., 2024a) 94.47±$\pm 1.37$ 95.10±$\pm 1.13$ 94.55±$\pm 0.86$ $94.98\pm 1.16$
Forest (Schimanski et al., 2024a) 95.37±$\pm 0.92$ 95.34±$\pm 0.94$ 94.78±$\pm 0.48$ 95.29±$\pm 0.65$
Biodiversity (Schimanski et al., 2024a) 92.76±$\pm 1.01$ 92.49±$\pm 1.03$ 92.46±$\pm 1.54$ 92.29±$\pm 1.23$
Nature (Schimanski et al., 2024a) 94.19±$\pm 0.81$ 93.50±$\pm 0.64$ 93.97±$\pm 0.26$ 93.55±$\pm 0.72$
(b) F1-scores of multiple models compared to EnvironmentalBERT (based on RoBERTa and distilRoBERTa).
@@
Multiple Benchmarks (Thulke et al., 2024) 71.4 77.1
(c) F1-scores of ClimateGPT-2 and custom performances (average Accuracy and F1-score over multiple benchmarks) of ClimateGPT compared to similarly sized models.
Table 1. Performance of domain-specific models on domain-specific tasks. The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: CL for Climate, Dis for Distil, EnvDisRob for EnvDistilRoBERTa, Evid. for evidences. Each row reports the performances of models fine-tuned in the same experimental setting. Detailed performances are reported in Appendix C.
As shown in Table 1(c), both ClimateGPT-2 and ClimateGPT significantly outperform the baseline model in their designated tasks. These models benefit from additional training on climate-related data, highlighting the advantages of domain adaptation. However, when trained from scratch, Thulke et al. (2024) report substantially lower performance, underscoring the importance of large-scale pretraining data. For smaller models (climateBERT, climateBUG-LM and EnvironmentalBERT), the domain-adaptation also enhanced the performance on domain-specific tasks, though the significance of this improvement is less pronounced. As detailed in Table 1(a), out of the 15 evaluations of fine-tuned domain-specific climateBERT (Webersinke et al., 2022) on climate-related tasks compared to the fine-tuned base model (distilRoBERTa), 4 reported an improvement compared to the base model (¿1%), 8 reported a marginal improvement (¡1%) and 3 indicate no measurable difference (¡=0%). When compared to DistilBERT in the last two tasks from Table 1(a), climateBERT exhibits a slight improvement in one case, while performing comparably in the other. Additionally, climateBERT achieves performance similar to the simple baselines in certain instances (Table 1(a)). For example, in Specificity Classification (Bingler et al., 2023), climateBERT outperforms an SVM with bag-of-word by 2%, and on the TCFD classification (Sampson et al., 2022) climateBERT performs slightly below the TF-IDF baseline model. Likewise, climateBUG-LM (Yu et al., 2024) demonstrates a slight improvement of less than 1% over BERT and FinBERT on a climate-related detection task. As shown in Table 1(c), the domain-adapted models shows a small improvement in the performance on the Environmental task, by 1.38% for distilRoBERTa and 0.84% for RoBERTa.
@@
Domain-specific language models, particularly those developed for environmental and climate-related texts, offer an interesting avenue for improving model performance through enhanced knowledge within specialized domains. Theoretically, such models should outperform general models by leveraging their tailored vocabulary and contextual understanding. However, our findings suggest that the improvements provided by the domain-specific models on the proposed domain-specific downstream tasks are limited. Indeed, we found that fine-tuning general models on domain-specific datasets consistently yields competitive performance. However, these findings do not undermine the significance of existing works that have been fundamental in demonstrating the potential of domain-specific models, particularly in generating domain-specific text (MLM and next token prediction). Moreover, ClimateGPT (Thulke et al., 2024) demonstrated significant improvement for the domain-adapted Llama-2, showing that domain-adaptation is highly relevant. Future research could focus on designing tasks where a lack of domain understanding—whether in terms of vocabulary or knowledge—significantly hinders performance.
4.2. Climate-Related Topic Detection
@@
ClimateBug-data (Yu et al., 2024) sentences from Banks’ reports relevant/ irrelevant Climate change and sustainability (including ESG, SDGs related to the environment, recycling and more)
ClimateBERT’s climate detection (Bingler et al., 2023) paragraphs from reports 1/0 Climate policy, climate change or an environmental topic
Climatext (Wikipedia, 10-K, claims) (Leippold and Varini, 2020) sentences from Wikipedia, 10-Ks or web scraping 1/0 Directly related to climate-change
Climatext (wiki-doc) (Leippold and Varini, 2020) sentences from a Wikipedia page 1/0 Extracted from a Wikipedia page related to climate-change
Sustainable signals’s reviews (Lin et al., 2023b) online product reviews (user comments) relevant/ irrelevant Contains terms related to sustainability
Table 2. Label definitions of the datasets related to climate change and sustainability topic detection.
Unsupervised Solutions
The earliest work on identifying corporate climate-related text was conducted by Doran and Quinn (2009), where the authors collected a large corpus of 10-K filings from 1995 to 2008, and filtered climate-related reports using hand-selected keywords. While Doran and Quinn (2009) focused on identifying climate-related text within company reports to assess how businesses addressed climate change, other early efforts took a different approach by analyzing the external impacts of media attention toward climate change on companies. For instance, Engle et al. (2020) developed the WSJ Climate Change News Index based on identifying climate-related news articles in media sources using climate-change vocabulary frequency. This index served as a tool to measure the influence of climate-change attention in the media on companies. Gorovaia and Makrominas (2025) proposed measuring the environmental content of CSR (corporate social responsibility) reports using a lexicon-based approach, with the goal of comparing CSR reports of environmental violators to companies with clean records.
@@
More recent studies have introduced several annotated datasets aimed at climate-related corporate text classification. One of the most prominent is ClimaText (Leippold and Varini, 2020), a large dataset from diverse sources such as Wikipedia, U.S. Security and Exchange Commission 10-K filings, and web-scraped content. The dataset is designed to cover a broad range of topics and document types to help assess climate-related discussions across different domains. It is divided into multiple subsets: wiki-doc, Wikipedia, 10-k, claims. Climatext (Wiki-doc) is annotated automatically, while Climatext (Wikipedia, 10-k and claims) is annotated by humans. In contrast, ClimateBERT’s authors (Webersinke et al., 2022) provide a smaller, more focused dataset consisting of paragraphs from companies’ annual and sustainability reports. This dataset was created as a downstream task for evaluating the ClimateBERT model, focusing on corporate disclosures. The dataset was later extended by Bingler et al. (2023), increasing its size and refining the annotation to capture both specificity and sentiment related to corporate climate discourse (see Sections 4.44.6). Another large dataset, ClimateBUG-data (Yu et al., 2024), also focuses on corporate communications but in EU banks. Beyond corporate disclosures, Sustainable Signals (Lin et al., 2023b) introduces a dataset of product reviews, each classified based on its relevance to sustainability. In addition to focusing on consumer perspectives through reviews, this dataset also examines sustainability aspects in product descriptions, providing a broader understanding of how sustainability is communicated in the context of consumer products.
@@
Supervised Solutions
The solutions proposed to tackle that task range from keyword-based models (Leippold and Varini, 2020; Bingler et al., 2023) to fine-tuning BERT-like models (Leippold and Varini, 2020; Webersinke et al., 2022; Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table LABEL:tab:models in the appendix for details). The performance of the best models is above 90% (Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table 12 in the appendix for details).

### v2

@@ 3.2 Definition of Greenwashing
As climate-related disclosures face increasing regulation and greenwashing poses significant risks to a company’s reputation, some companies adopt a strategy of silence, avoiding discussions about their environmental impact. This deliberate lack of communication is known as greenhushing [Letzing]. These definitions imply that greenwashing is a deliberate act; however, in many cases, it results from an error or miscommunication by companies genuinely trying to showcase their sustainability efforts, and that are trying to best follow disclosure standards.
As shown by Definitions 1 and 2, greenwashing is not defined by easily identifiable properties but as a general concept. Since the concept is so unspecific, researchers focused on components indicative of potentially misleading communications but easier to define. Because of this, in this review we are mentioning “greenwashing” explicitly, but also paraphrasing it as “misleading communications”, “misrepresentation of the company’s environmental impact/stance/performance”, or mentioning only components of it such as “cheap talk”, “selective disclosure/transparency”, “deceptive techniques”, “biased narrative”. They should all be understood in the context of climate-related misleading communications, as components associated or indicative of potential greenwashing even if they are not synonymous.
4 Foundational Tasks for Greenwashing Detection
4.1 Pretraining Models on Climate-Related Text
The first step in applying a language model for a given task is typically the pretraining process, which involves training the model on relevant corpora. Although general-purpose models such as BERT [devlin-etal-2019-bert] and LLaMA [touvron2023llamaopenefficientfoundation] have demonstrated strong performance in various tasks, they are inherently limited by the knowledge and vocabulary present in their training corpora. This has motivated the development of climate-focused language models, which aim to encode domain knowledge directly into the pretrained representations.
@@ 4.1 Pretraining Models on Climate-Related Text
A language model is pretrained on climate-related text, to produce a domain-specific language model that will be subsequently fine-tuned to specific tasks. There are two approaches:
  1. (i)
    Training a model with a domain-specific corpus from scratch, or
  2. (ii)
    Further-training [^2] Further-training refers to the process of training a pretrained model on its original task using additional domain-specific data to specialize its knowledge for that domain. a generalist pretrained model on a domain-specific corpus.
@@
A variety of climate-related corpora have been assembled for this purpose, ranging from scientific abstracts and policy documents to sustainability reports and ESG disclosures. These corpora differ in scope and quality, and only a few are publicly accessible. The lack of standardized, openly available pretraining datasets remains a barrier to consistent benchmarking and reproducibility. nicolas_webersinke_climatebert_2021 gathered climate-related news articles, research abstracts, and corporate reports. vaghefi2022deep introduced Deep Climate Change, a dataset composed of abstracts of articles from climate scientists, and built a corpus specific to climate research texts. schimanski_bridging_2023 introduced a dataset that focuses on text related to Environment Social and Governance (ESG). Similarly, Mehra_2022 built a dataset using text from the Knowledge Hub of Accounting for Sustainability for an ESG domain-specific corpus. More recently, thulke2024climategpt proposed a dataset comprising news, publications (abstracts and articles), books, patents, the English Wikipedia, policy and finance-related texts, Environmental Protection Agency documents, and ESG, and IPCC reports. mullappilly-etal-2023-arabic proposed a climate-specific multilingual dataset. yu_climatebug_2024 published a pretraining dataset constructed with annual and sustainable reports from EU banks. Unfortunately, to the best of our knowledge, only [yu_climatebug_2024]’s ClimateBUG-data dataset is publicly available.
Models
Based on the above datasets, the models ClimateBERT, ClimateGPT-2, climateBUG-LM and EnvironmentalBERT, ESGBERT were proposed [nicolas_webersinke_climatebert_2021, vaghefi2022deep, yu_climatebug_2024, schimanski_bridging_2023, Mehra_2022]. More recently, generative domain-specific models such as Llama-2 for ClimateGPT [thulke2024climategpt] or Vicuna for Arabic Mini-ClimateGPT [mullappilly-etal-2023-arabic] have also been proposed.
@@
The domain-specific models drastically improve the performance on domain-specific masked-language modeling [nicolas_webersinke_climatebert_2021, yu_climatebug_2024] and next token prediction [vaghefi2022deep]. They were also evaluated on domain-specific downstream tasks. These tasks are either based on pre-existing datasets such as ClimateFEVER [diggelmann_climate-fever_2020] or introduced by the authors, as in ClimateBERT’s climate detection [nicolas_webersinke_climatebert_2021]. At this stage, we evaluate whether fine-tuning enhances downstream performance, with a detailed analysis of the tasks presented in subsequent sections.
Table 1: Performance of ClimateBERT on domain-specific tasks (using F1-score, except for TCFD classification which is using ROC-AUC). The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: CL for Climate, Dis for Distil, Evid. for evidences. Each row reports the performances of models fine-tuned in the same experimental setting. Detailed performances are reported in Appendix C.
Cl. BERT Dis. RoBERTa NB
Climatext [spokoyny2023answering, varini_climatext_2020] 85.14 86.06 83.39
[…]
Cl. BERT Dis. RoBERTa TF-IDF
TCFD classification [sampson_tcfd-nlp_nodate] 0.852 0.819 0.867
Cl. BERT Dis. RoBERTa SVM
SciDCC [spokoyny2023answering, mishra2021neuralnere] 52.97 51.13 48.02
ClimaTOPIC [spokoyny2023answering] 64.24 63.61 58.34
Cl.FEVER (claim) [xiang_dare_2023, diggelmann_climate-fever_2020] 76.8 72
[…]
Cl.FEVER (evid.) [spokoyny2023answering, diggelmann_climate-fever_2020] 61.54 61.54
Cl.Stance [spokoyny2023answering, vaid-etal-2022-towards] 52.84 52.51 42.92
ClimateEng [spokoyny2023answering, vaid-etal-2022-towards] 71.83 72.33 51.81
ClimaINS [spokoyny2023answering] 84.80 84.38 86.00
ClimaBENCH [spokoyny2023answering] 69.44 69.27
Nature [Schimanski2024nature] 94.11 94.03
@@
Commitment&Actions [bingler2023cheaptalkspecificitysentiment] 81 79 76
Specificity [bingler2023cheaptalkspecificitysentiment] 77 76 75
Cl. BERT Dis. BERT SVM
Env. Claim [stammbach_environmental_2023] 83.8 83.7 70.9
Cl. BERT Dis. BERT LSTM
DARE sentiment [xiang_dare_2023] 87.4 89.9 88.2
Table 2: F1-scores of multiple models compared to EnvironmentalBERT. The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: Dis for Distil, EnvDisRob for EnvDistilRoBERTa. Each row reports the performances of models fine-tuned in the same experimental setting. Detailed performances are reported in Appendix C.
@@
[…]
@@
[…]
Table 3: F1-scores of ClimateGPT-2 and custom performances of ClimateGPT compared to similarly sized models on domain-specific benchmarks. The figure displayed in the tables are the values reported by authors in the corresponding studies. Detailed performances are reported in Appendix C.
@@
Multiple Benchmarks [thulke2024climategpt] 71.4 77.1
As shown in Table 1 and Table 2, evaluations on climate-related downstream tasks show mixed but informative results: (i) domain-adapted models often outperform their base counterparts on specialized tasks, (ii) gains are frequently modest, and (iii) strong generalist baselines—when fine-tuned on the same domain-specific datasets—remain highly competitive. These findings indicate that domain adaptation helps encode environmental terminology and discourse structure, but its benefits are task-dependent and not uniformly substantial. In Table 3, both ClimateGPT-2 and ClimateGPT significantly outperform the baseline model. These models benefit from additional training on climate-related data, highlighting the advantages of domain adaptation. However, when trained from scratch, thulke2024climategpt report substantially lower performance, underscoring the importance of large-scale pretraining data.
@@
In theory, domain-specific language models should outperform general models by leveraging their tailored vocabulary and contextual understanding. However, our findings suggest that the improvements provided by these models on the proposed domain-specific downstream tasks are limited. Indeed, we found that fine-tuning general models on domain-specific datasets consistently yields competitive performance. These findings do not undermine the significance of existing works that have been fundamental in demonstrating the potential of domain-specific models, particularly in generating domain-specific text. Moreover, ClimateGPT [thulke2024climategpt] demonstrated significant improvement for the domain-adapted Llama-2, showing that domain-adaptation is highly relevant. Future research could focus on designing tasks where a lack of domain understanding—whether in terms of vocabulary or knowledge—significantly hinders performance.
4.2 Climate-Related Topic Detection
@@
Table 4: Label definitions of the datasets related to climate change and sustainability topic detection.
@@
ClimateBug-data yu_climatebug_2024 Sentences from bank reports relevant/ irrelevant Climate change and sustainability (including ESG, SDGs related to the environment, recycling and more)
ClimateBERT’s climate detection bingler2023cheaptalkspecificitysentiment Paragraphs from reports 1/0 Climate policy, climate change or an environmental topic
Climatext (Wikipedia, 10-K, claims) varini_climatext_2020 Sentences from Wikipedia, 10-Ks or web scraping 1/0 Directly related to climate-change
Climatext (wiki-doc) varini_climatext_2020 Sentences from a Wikipedia page 1/0 Extracted from a Wikipedia page related to climate-change
Sustainable signals’s reviews linSUSTAINABLESIGNALSijcai2023 Online product reviews (user comments) relevant/ irrelevant Contains terms related to sustainability
Annotation-free
The earliest work on identifying corporate climate-related text was conducted by doran_risk_disclosure, where the authors collected a large corpus of 10-K filings from 1995 to 2008, and filtered climate-related reports using hand-selected keywords. While doran_risk_disclosure focused on identifying climate-related text within company reports to assess how businesses addressed climate change, other early efforts took a different approach by analyzing the external impacts of media attention toward climate change on companies. For instance, Engle_hedging_climatechange developed the WSJ Climate Change News Index based on identifying climate-related news articles in media sources using climate-change vocabulary frequency. This index served as a tool to measure the influence of climate-change attention in the media on companies. csr_report_greenwashing proposed measuring the environmental content of CSR (corporate social responsibility) reports using a lexicon-based approach, with the goal of comparing CSR reports of environmental violators to companies with clean records.
@@
More recent studies have introduced several annotated datasets aimed at climate-related corporate text classification. One of the most prominent is ClimaText [varini_climatext_2020], a large dataset from diverse sources such as Wikipedia, U.S. Security and Exchange Commission 10-K filings, and web-scraped content. The dataset is designed to cover a broad range of topics and document types to help assess climate-related discussions across different domains. It is divided into multiple subsets: wiki-doc, Wikipedia, 10-k, claims. Climatext (Wiki-doc) is annotated automatically, while Climatext (Wikipedia, 10-k and claims) is annotated by humans. In contrast, ClimateBERT’s authors [nicolas_webersinke_climatebert_2021] provide a smaller, more focused dataset consisting of paragraphs from companies’ annual and sustainability reports. This dataset was created as a downstream task for evaluating the ClimateBERT model, focusing on corporate disclosures. The dataset was later extended by bingler2023cheaptalkspecificitysentiment, increasing its size and refining the annotation to capture both specificity and sentiment related to corporate climate discourse (see Sections 4.44.6). Another large dataset, ClimateBUG-data [yu_climatebug_2024], also focuses on corporate communications but in EU banks.
Beyond corporate disclosures, Sustainable Signals [linSUSTAINABLESIGNALSijcai2023] introduces a dataset of product reviews, each classified based on its relevance to sustainability. In addition to focusing on consumer perspectives through reviews, this dataset also examines sustainability aspects in product descriptions, providing a broader understanding of how sustainability is communicated in the context of consumer products.
@@
Models
The solutions proposed to tackle that task range from keyword-based models [varini_climatext_2020, bingler2023cheaptalkspecificitysentiment] to fine-tuning BERT-like models [varini_climatext_2020, nicolas_webersinke_climatebert_2021, garridomerchán2023finetuning, bingler2023cheaptalkspecificitysentiment, yu_climatebug_2024]. The performance of the best models is above 90% [garridomerchán2023finetuning, bingler2023cheaptalkspecificitysentiment, yu_climatebug_2024]. [^3] See Table 13 in the appendix for details.